OpenStudy (anonymous):

Multiply and simplify: (3 + √3) (5 - 4√3) Answers: 3 15-4√3 3-7√3

OpenStudy (anonymous):

Do you know how to do the problem? Try?

OpenStudy (anonymous):

Foil or whatever.

OpenStudy (owlcoffee):

Let's apply what is called "distributive": \[(3+\sqrt{3})(5-4\sqrt{3})\] we'll end up with: \[(3)(5)-(3)(4\sqrt{3})+(\sqrt{3})(5)-(\sqrt{3})(4\sqrt{3})\] think you can take over from here?

OpenStudy (anonymous):

ummm I don't really get it :/

OpenStudy (owlcoffee):

All I did was multiplying, he rest is pretty much using properties and simplifying.

OpenStudy (anonymous):

okay, ill try

OpenStudy (anonymous):

You gotta learn to multiply.

OpenStudy (anonymous):

im lost, this is helpless -.-

OpenStudy (anonymous):

i don;t even know where to start...

OpenStudy (owlcoffee):

Oh Wio, don'te be so mean :(. Charlie, start by operating the most simple thing.

OpenStudy (anonymous):

that the point, I don't know where that even is

OpenStudy (anonymous):

I don't mean learn to multiply like arithmetic, I mean learn to multiply binomials.

OpenStudy (anonymous):

3-7sqrt(3)